OpenStudy (anonymous):

decker has one allot containing 27% copper and another containing 60% copper. How many pounds of each must he use to make 49 pounds of a third alloy containing 39% copper? Round to the nearest hundredth.

OpenStudy (anonymous):

0.27x+0.6y=49(0.39) x+y=49 That's your system of equations and you can solve from there.